1.   A standing wave has a frequency of 324 HZ.  It is vibrating in 4 sections.  Calculate the velocity of the wave if the length is 3 metres.  L=n λ/2   (/2)

  

 

 

 

B.  If the length of the string was shortened to 2 metres, determine the period of the wave?  (/2)

 

 

 C.  Determine the fundamental frequency (n=1) for both lengths of string.  (/2)




1
Expert's answer
2021-06-21T15:06:46-0400

Gives

Frequency (n)=324Hz

Section=4

We know that

"L=\frac{n\lambda}{2}"

"L=3m"

"\nu=324Hz"

n=3

"\lambda=\frac{2\times3}{3}=2m"

"v=\nu\lambda"

Put value

"v=324\times2=648m/sec"

Part(2)

Length(l)=2m


"v=\frac{\lambda}{T}"

"T=\frac{\lambda}{v}"

"L=\frac{n\lambda}{2}"

"2=\frac{1 \times\lambda}{2}"

"\lambda=4m"

"T=\frac{4}{648}=0.0061sec"

Part(3)

When n=1 ,"\lambda=6m"

"\nu=\frac{v}{\lambda}=\frac{648}{6}=108Hz"

n=1,"\lambda=4m" ",l=2m"

"\nu=\frac{v}{\lambda}"

V=491m/sec

"\nu=\frac{491}{2}=245Hz"